In a right triangle, the hypotenuse has a length of 20cm. One of the angles in this triangle is five times another angle(in degrees). solve the triangle.
Answer by LinnW(1048)

x first angle
y second angle
y = 5x
x^2 + y^2 = 20^2
x^2 + (5x)^2 = 20^2
x^2 + 25x^2 = 400
26x^2 = 400
divide each side by 26
x^2=400/26
take square root of each side
x = 3.92
y=5(3.92)
y=19.6